please help I can't figure this out
2x-y=-5
-2x -2x
-y=-5-2x
-1 -1 -1
y=5+2x
2x-y=-5
2()-()=-5

Answer by sdmmadam@yahoo.com(530)

please help I can't figure this out
2x-y=-5
-2x -2x
-y=-5-2x
-1 -1 -1
y=5+2x
2x-y=-5
2()-()=-5

But what is the need to do all this?
May be you are askedto present this given linear equation in x and y
in the slope and y-intercept form! Even that can be done by simply transferring (-y) from the left to the right and bringing (-5) from the right to the left.
Let me show you what you have done.
2x-y=-5
Adding (-2x) from the left on both the sides
(-2x) + (2x-y) = (-2x) -5
[(-2x)+(2x)]-y = -2x-5
0-y = -2x-5
Multiplying by (-1)
(-1)X(-y)= (-1)X(-2)x+(-1)X(-5)
y = 2x+5
slope = 2 and y-intercept = 5